Is this correct?
Solve for x:
3ln(2x+3)-2=1
3ln(2x+3)/3=3/3
ln(2x+3)=1
2x+3=e
2x/2=-.2817/2
x= -.1409

Answer by jsmallt9(3296)

From here:
2x+3=e
I would finish as follows:
2x = e-3
x = (e-3)/2

Unless I needed or was asked to find a decimal approximation for the answer, I would not replace e with a decimal approximation for it, as you did,